# Swiss Army Knife: Synergizing BiAses in Knowledge from Vision Foundation Models for Multi-Task Learning

### v1

@@ B.1 Models
Architecture. We use DINOv2 (Oquab et al., 2024), CLIP (Radford et al., 2021), and SAM (Kirillov et al., 2023) as our VFM teachers, with pretrained models from timm (Wightman, 2019). For CLIP, we utilize the OpenCLIP (Cherti et al., 2023) reproduction in our experiments. The model names of the VFMs and multi-teacher VFM distillation baselines, RADIO (Ranzinger et al., 2024) [^1] https://github.com/NVlabs/RADIO and Theia (Shang et al., 2024) [^2] https://huggingface.co/collections/theaiinstitute/theia-66a7a6ae80a707547c358cce, are listed in Table 5. For the Teacher-Agnostic Stem in the SAK student model, we employ ViT-S, ViT-B, and ViT-L backbones, initialized with pretrained weights from timm. While we keep the CLS tokens in the model, we only use patch tokens as representations for both the teachers and student models. Other modules in the student are trained from scratch. The adapters in the Teacher-Specific Adapter Path use a channel reduction ratio of 4.
@@
Input resolution. As detailed in Table 1, the VFMs are pretrained on diverse image resolutions. Additionally, RADIO-B is pretrained on $768\times 768$, RADIO-L is pretrained on $1024\times 1024$, and Theia on $224\times 224$. In our experiments, we use a unified input size for all models, except SAM. In the first stage, we use an input size of $384\times 384$, while in the second stage, the input size is set to $512\times 512$ for PASCAL-Context and $448\times 576$ for NYUD-v2. For SAM, we pad and resize all images to $1024\times 1024$. Feature maps from all models are interpolated to $1/4$ of the input resolution before being decoded by the heads for a fair comparison.
@@ B.6 Training
As outlined in Section 3.4, the first stage trains the student model, which consists of the TAS and TSAP modules, with the distillation loss. Then in the second stage, we add the MoR Routers and prediction heads, training the entire student model with a combination of distillation loss and task-specific losses. The loss balancing factor $\gamma$ in Equation 6 is set as $1.0$ for simplicity. Note that VFM teachers are always frozen in both stages. The detailed hyperparameter configurations are listed in Table 6, and the second stage follows the setup used in prior work (Ye & Xu, 2023b). We implement our methodology with PyTorch (Paszke et al., 2019) and run experiments on NVIDIA H100 GPUs with 96G VRAM.
@@ Appendix C Additional Experimental Results
Appendix D Additional Related Work
D.1 Knowledge Distillation
@@
Knowledge Distillation (KD) is a model compression technique where a smaller student model learns from a larger teacher model, typically using the teacher’s output logits as soft targets to guide training (Buciluǎ et al., 2006; Hinton et al., 2014; Ba & Caruana, 2014; Beyer et al., 2022). To improve knowledge transfer, some studies have proposed using intermediate representations (Romero et al., 2015; Ahn et al., 2019; Heo et al., 2019; Zagoruyko & Komodakis, 2017). Multi-teacher knowledge distillation has also been explored, where knowledge is transferred from multiple teacher models simultaneously or progressively to improve the student’s generalization (Sau & Balasubramanian, 2016; You et al., 2017; Fukuda et al., 2017; Wen et al., 2024; Cao et al., 2023). Another branch of work, known as knowledge amalgamation, combines knowledge from multiple teachers tackling different tasks into a student that learns the union of all tasks (Shen et al., 2019a; b; Luo et al., 2019; Ye et al., 2019b; Vongkulbhisal et al., 2019; Thadajarassiri et al., 2021; 2023). Such methods mainly unify classification tasks from multiple domains. In contrast, our approach addresses multiple heterogeneous vision tasks (e.g., semantic segmentation, surface normal estimation) simultaneously.
D.2 Vision Foundation Models

### v2

@@ B.1 Models
Architecture. We use DINOv2 (Oquab et al., 2024), CLIP (Radford et al., 2021), and SAM (Kirillov et al., 2023) as our VFM teachers, with pretrained models from timm (Wightman, 2019). For CLIP, we utilize the OpenCLIP (Cherti et al., 2023) reproduction in our experiments. The model names of the VFMs and multi-teacher VFM distillation baselines, RADIO (Ranzinger et al., 2024b) [^1] https://github.com/NVlabs/RADIO and Theia (Shang et al., 2024) [^2] https://huggingface.co/collections/theaiinstitute/theia-66a7a6ae80a707547c358cce, are listed in Table 5. For the Teacher-Agnostic Stem in the SAK student model, we employ ViT-S, ViT-B, and ViT-L backbones, initialized with pretrained weights from timm. While we keep the CLS tokens in the model, we only use patch tokens as representations for both the teachers and student models. Other modules in the student are trained from scratch. The adapters in the Teacher-Specific Adapter Path use a channel reduction ratio of 4.
@@
Input resolution. As detailed in Table 1, the VFMs are pretrained on diverse image resolutions. Additionally, RADIO-B is pretrained on $768\times 768$, RADIO-L is pretrained on $1024\times 1024$, and Theia on $224\times 224$. In our experiments, we use a unified input size for all models, except SAM. In the first stage, we use an input size of $384\times 384$, while in the second stage, the input size is set to $512\times 512$ for PASCAL-Context and $448\times 576$ for NYUD-v2. For SAM, we pad and resize all images to $1024\times 1024$. Feature maps from all models are interpolated to $1/4$ of the output resolution before being decoded by the heads for a fair comparison.
@@ B.6 Training
As outlined in Section 3.4, the first stage trains the student model, which consists of the TAS and TSAP modules, with the distillation loss. Then in the second stage, we add the MoR Routers and prediction heads, training the entire student model with a combination of distillation loss and task-specific losses. The loss balancing factor $\gamma$ in Equation 6 is set as $1.0$ for simplicity. Note that VFM teachers are always frozen in both stages. The detailed hyperparameter configurations are listed in Table 6. For the first stage, we follow the setup of RADIO (Ranzinger et al., 2024b) for optimizer, learning rate, LR scheduler, and distillation loss in Equation 5, while adopting Theia’s (Shang et al., 2024) batch size configuration since we use the same dataset. To enhance training efficiency, we reduce the number of epochs. For the second stage follows the identical hyperparameters used in prior multi-task learning works (Ye & Xu, 2022; 2023b; Yang et al., 2024d). We implement our methodology with PyTorch (Paszke et al., 2019) and run experiments on NVIDIA H100 GPUs with 96G VRAM.
@@ Appendix C Additional Experimental Results
Upper-bound results of teacher amalgamation. To evaluate the upper-bound performance of amalgamating the VFM teachers, we build an upper-bound model using the frozen encoders of three VFM teachers—DINOv2, CLIP, and SAM—along with a learnable ViT encoder as a surrogate for the Teacher-Agnostic Stem (TAS). All components employ a ViT-B backbone. For fusing the representations, we explore three strategies: element-wise addition, channel concatenation, and our proposed Mixture-of-Representations (MoR).
Table 14: Comparison with upper-bound models on PASCAL-Context, based on ViT-B backbones.
Model Fuse #Param MACs Semseg Parsing Saliency Normal Boundary ${\Delta_{m}\%\uparrow}$
[…]
Upper-bound Addition 378M 1091G 80.27 71.44 84.74 13.82 72.69 -0.47
Concat 820M 7661G 82.26 73.85 84.63 13.98 75.55 1.21
MoR 384M 1097G 80.58 72.73 84.88 14.05 74.01 0.02
[…]
Table 14 shows that a naive addition of representations fails to fully leverage the teachers’ knowledge, as mentioned in the introduction and further supported by our ablation study in Section 4.3 and Table 3. While channel concatenation mitigates this limitation and leads to upper-bound performance, it comes at the cost of a significant increase in parameter count and computational overhead due to the expanded dimensionality of the fused representations. In contrast, our MoR approach surpasses addition by 0.5% while maintaining comparable parameters and MACs, thereby demonstrating its effectiveness even in this alternative setup.
Compared to these upper-bound models, our SAK outperforms the naive addition with around 1/3 of the parameters and half of the computational cost. This further validates the effectiveness and efficiency of our method.
@@
Appendix D Additional Discussions
D.1 Model Efficiency
Training cost. Since SAK employs a multi-teacher distillation framework, the iterative forwarding of VFM teachers inevitably increases training costs, which cannot be overlooked. However, this process is a common requirement for all multi-teacher distillation methods, and when compared to other baselines, such as RADIO, SAK demonstrates remarkable efficiency.
In Table D.1, we calculate the number of parameters and computational costs for teachers and students in RADIO and SAK, both using a ViT-L backbone for the student. The results show that RADIO incurs 2.5 times the parameters and forwarding costs of SAK. Additionally, RADIO’s larger input size of 1,024 introduces more tokens in the representations, further reducing the backward efficiency. Hence, SAK shows significantly lower demands in RAM, GPU memory, training time, and storage.
Table 16: Comparison with RADIO on number of parameters and computational cost (measured by MACs, Multiply–accumulate operations) of all teachers and students. The calculation is based on students with ViT-L backbones for both methods. SAK uses VFM teachers with ViT-L backbones.
RADIO teachers and students Input size #Param MACs SAK teachers and students Input size #Param MACs
DFN CLIP-H/14 378 631M 460G DINOv2-L/14 384 304M 221G
SigLIP-SO400M/14 384 413M 300G OpenCLIP-L/14 384 304M 221G
DINOv2-g/14 224 1135M 291G SAM-L/16 1024 307M 1308G
SAM-H/16 1024 636M 2730G
RADIOv2.5-L/16 1024 320M 1240G SAK-L/16 384 343M 198G
[…]
Moreover, as listed in Table 24, RADIO is trained on the large-scale DataComp-1B dataset (Gadre et al., 2023) (1.4B images) with 614M total samples seen. In contrast, SAK is trained on ImageNet-1k dataset (1.2M images) with only 36M samples seen, further highlighting the efficiency of SAK in training and memory costs.
Inference cost. It is important to note that our framework does not require teachers during inference. Instead, the representations distilled from the teachers are generated directly by the lightweight TSAP modules, which introduce minimal additional parameters and computational overhead. To provide a clearer perspective, we present the number of parameters and computational cost introduced by our modules when integrated into the ViT-B or ViT-L backbone. Table 17 indicates that each TSAP branch accounts for less than 5% of the backbone’s parameters and computations (4M vs. 86M parameters, 4G vs. 88G MACs with ViT-B). Similarly, each MoR Router is as lightweight as 2M parameters and 2G MACs, even with the larger ViT-L backbone.
Table 17: Number of parameters and computational cost when integrating different components into SAK, based on ViT-B or ViT-L backbone. The calculation is based on a student corresponding to three teachers on the five-task PASCAL-Context dataset.
Model #Param MACs Model #Param MACs
ViT-B/16 86M 88G ViT-L/16 304M 311G
+TSAP (3 branches) 98M 100G +TSAP (3 branches) 344M 351G
+MoR (5 tasks) 104M 106G +MoR (5 tasks) 354M 362G
D.2 Differences between MoE and MoR
Unlike Mixture-of-Experts (MoE) used in previous multi-task learning (Yang et al., 2024d; Chen et al., 2023; Liang et al., 2022; Ye & Xu, 2023a) or general-purpose models (Shazeer et al., 2017; Jacobs et al., 1991), where the experts are typically homogeneous during optimization, the experts in our Mixture-of-Representation (MoR) approach consist of representations distilled from various VFM teachers as well as the TAS. These experts inherently possess different knowledge, making them heterogeneous by design. As a result, the MoR Router does not require a balanced routing because task-specific routers dynamically weigh and combine the most relevant representations for each task. Empirically, we do not encounter any difficulties in training the MoR Routers. This approach ensures that the router effectively synergizes the complementary strengths of the diverse representations to optimize performance for different tasks.
D.3 VFM Teacher Selection
Principles of VFM teacher selection. It is well-established that DINOv2, CLIP, and SAM are among the most widely used Vision Foundation Models. Prior baselines RADIO and Theia also employ them for experiments. Besides, RADIO uses DFN CLIP (Fang et al., 2024) and SigLIP (Zhai et al., 2023), both of which are improved models derived from CLIP. However, as shown in the Table 18, DFN CLIP and SigLIP fail to attain comparable outcomes to CLIP. For this reason, we choose to adhere to CLIP when selecting teachers.
Table 18: Comparison of CLIP, DFN CLIP, and SigLIP on PASCAL-Context, based on ViT-B backbones.
Model Fine-tune Semseg Parsing Saliency Normal Boundary ${\Delta_{m}\%\uparrow}$
[…]
CLIP (Radford et al., 2021) MT Full 76.84 65.87 84.61 13.91 70.50 -3.66
Freeze 78.33 65.31 81.48 16.84 67.43 -9.33
[…]
Freeze 69.35 53.48 81.36 17.24 64.85 -16.24
[…]
Freeze 51.72 45.34 77.93 18.35 60.33 -26.61
Theia utilizes ViT (pretrained on ImageNet) and Depth Anything (Yang et al., 2024b) as additional teachers. Since the Teacher-Agnostic Stem in SAK is initialized with an ImageNet-pretrained ViT backbone, we do not include ViT as an additional teacher to reduce training costs. While Depth Anything is a high-impact VFM, we leave its exploration for future work.
Additionally, we avoid using very large models, e.g., ViT-H or ViT-g based, in our experiments due to constraints in computational resources (RADIO uses 64 GPUs, while we can only use 8 GPUs). It is worth noting that SAK is a highly flexible framework, capable of integrating and benefiting further from more powerful VFM teachers, making it adaptable to various settings.
Sensitivity of SAK to VFM teacher selection. We supplement another analysis with a SAK student distilled from DINOv2 and SigLIP teachers, with ViT-B backbones for both teachers and student. As shown in Table 19, compared with the student distilled from DINOv2 and CLIP teachers (presented in Table 4), it exhibits better performance on three out of five tasks, while achieving comparable overall accuracy, further underscoring the robustness and adaptability of our method.
Table 19: Performance w.r.t. different combinations of VFM teachers on PASCAL-Context.
Teachers Semseg Parsing Saliency Normal Boundary ${\Delta_{m}\%\uparrow}$
[…]
DINOv2+CLIP 81.53 71.95 84.49 14.16 72.59 -0.60
DINOv2+SigLIP 81.41 71.03 84.62 14.06 72.92 -0.63
D.4 Framework Extension
As highlighted in our introduction, SAK is a highly flexible framework that can be easily adjusted to add new teachers or new downstream tasks.
Adding new VFM teachers. When adding a new teacher, we freeze the existing TAS and TSAP modules to preserve the knowledge transferred from the previous teachers A new TSAP module is then introduced and distilled for the new teacher, facilitating efficient adaptation. In the second stage, we train on the downstream dataset with all VFM teachers and tune the entire student model.
We conduct an experiment by incorporating a new SigLIP teacher into a student already distilled from DINOv2, CLIP, and SAM. As shown in Table 20, including SigLIP teacher boosts performance on Semseg and Normal tasks and yields competitive results on Saliency and Boundary tasks. However, since SigLIP encodes knowledge similar to CLIP during pretraining and does not surpass CLIP on the benchmark, it is reasonable that it could not lead to further improvements. Moreover, the additional representations may increase the difficulty for the MoR Routers in learning optimal weights, potentially accounting for the degradation in human parsing.
Table 20: Performance of SAK with additional SigLIP teacher on PASCAL-Context.
Teachers Semseg Parsing Saliency Normal Boundary ${\Delta_{m}\%\uparrow}$
[…]
DINOv2+CLIP+SAM 81.65 72.38 84.87 14.05 73.23 -0.03
+SigLIP 81.78 70.66 84.85 14.02 73.21 -0.45
Adding different tasks. Additionally, when applying to different downstream tasks, it is only necessary to add task-specific decoders and perform the second-stage training (with distillation) or fine-tuning (without distillation), as the first-stage distillation is agnostic to the downstream tasks. This process is significantly more efficient than the first stage. Meanwhile, we can readily remove a branch of TSAP if it is no longer necessary, without affecting the functionality of the TAS or other TSAP branches, ensuring the modularity and adaptability of our design. We have also shown in Figure 5 and its accompanying analysis that SAK exhibits strong robustness and generalization capability in downstream tasks.
D.5 Comparison of SAK with Teachers
Why cannot SAK outperform teachers at every task? Firstly, performance loss is inevitable during distillation, since our distillation process is lightweight compared to the extensive pretraining of VFM teachers. As detailed in Table 1, our VFM teachers are pretrained on large-scale datasets containing hundreds to thousands of times more image samples than the ImageNet-1k dataset used for distillation. Moreover, the Teacher-Specific Adapter Path, designed to preserve teachers’ biases, utilizes a standard and simple architecture with less than 5% of the backbone’s parameters and computations. Consequently, it is reasonable that the distilled SAK cannot fully inherit teachers’ knowledge when compressing three teacher models into a single student of similar model size, accounting for why our approach cannot outperform the teachers on their proficient tasks. We believe that leveraging larger datasets could further enhance knowledge transfer during distillation and benefit downstream tasks. Another reason is that SAM operates at a higher input resolution of 1,024, rather than 512 for other models, which bolsters its strength in tasks like boundary detection that rely heavily on fine image details.
Why can SAK outperform teachers on specific tasks? The observed superiority of SAK over teachers stems from our motivation to synergize the complementary strengths of teachers. Complementarity exists not only across tasks but also within tasks. Take semantic segmentation as an example, as illustrated in Figure 1 and qualitative analysis in Section 2.2, DINOv2 excels at capturing localized features but CLIP offers strong object-level understanding with its rich semantic knowledge from the language domain. SAM produces exceptional fine-grained pixel-level masks due to its higher resolution. Additional analysis of representations in Appendix A further supports these observations. By preserving the intra-task representation biases during distillation and amalgamating them using proposed Mixture-of-Representations, SAK achieves improved performance.
Another contributing factor is the balance between common knowledge and task-specific information. Regarding saliency estimation and surface normal estimation, three VFM teachers perform suboptimally when frozen, as they lack downstream task-specific information. Though fine-tuning can alleviate this limitation, it results in degradation of accuracy in segmentation tasks for DINOv2 and CLIP, as shown in the Table 7. This trade-off, known as the negative transfer problem in multi-task learning (Vandenhende et al., 2021; 2020), highlights the challenge of balancing task-specific and pretrained knowledge.
In contrast, SAK preserves the representation biases of the frozen teachers within the TSAP modules while leveraging the TAS to learn shared knowledge and downstream information. This disentanglement ensures knowledge diversity without mutual interference. Moreover, MoR Routers dynamically weigh and combine the most relevant representations for each task, bridging the gap between general-purpose knowledge and task-specific characteristics.
D.6 Evaluation on Additional Tasks
First, we emphasize that our work is less targeted at single-task learning. Instead, our key motivation lies in the inherent representation biases of VFMs, which result in advantages and disadvantages across different vision tasks. Our SAK synergizes these complementary biases to enhance multi-task learning. Therefore, the tasks we evaluated are defined by the datasets, since each image sample must be labeled for every task included. Nevertheless, we also provide additional evaluations on single-task learning, specifically for classification, depth estimation, and vision-language learning, to offer a more comprehensive analysis.
Linear classification. Following the setup of DINOv2, we evaluate linear probing on the ImageNet-1k dataset. We freeze the backbones and train a linear classifier and the MoR Routers of SAK using the same hyperparameters as DINOv2. For comparison, we also evaluate the VFM teachers and RADIO, while reporting Theia’s result from its original paper.
Table 21 shows that SAK outperforms the two baselines, despite a performance gap compared to the DINOv2 and CLIP teachers due to inevitable loss during distillation. It is worth noting that RADIO uses more powerful teachers, including DINOv2-g/14-reg (accuracy 87.1), while Theia uses DINOv2-L/14 teacher (accuracy 86.3).
Table 21: Comparison with VFMs and baselines on linear classification on ImageNet-1k, reported by Top-1 accuracy on the validation set. We freeze the backbone and perform linear probing at resolution of 224.
[…]
DINOv2 (Oquab et al., 2024) ViT-B/14 84.5
CLIP (Radford et al., 2021) ViT-B/16 84.7
SAM (Kirillov et al., 2023) ViT-B/16 46.9
RADIO (Ranzinger et al., 2024b) ViT-B/16 78.2
Theia (Shang et al., 2024) ViT-B/16 75.2
[…]
Depth estimation. We also evaluate the monocular depth estimation task on the NYUd dataset (Silberman et al., 2012) (note it is different from the NYUDv2 dataset used in our primary experiments), which is the proficient task of DINOv2. We follow the evaluation protocol (Li et al., 2024b) used by DINOv2 and use the identical experimental setups and decode heads (lin. 1 and lin. 4). Since DINOv2 does not provide a complete codebase for this task, we reproduce the pipeline. We also evaluate ViT, SAM, Theia, and include baselines reported in the DINOv2 paper, namely OpenCLIP (Cherti et al., 2023), MAE (He et al., 2022), and DINO (Caron et al., 2021).
As demonstrated in Table 22, while SAK does not fully match the teacher’s performance due to inevitable loss during distillation and different patch size, it surpasses other foundation models and baselines, even those with larger backbones. Moreover, the effectiveness of synergizing multiple VFM teachers is clearly evidenced by the results, as distillation from three teachers improves SAK’s performance, bringing it closer to the upper bound of the teacher.
Table 22: Comparison with VFMs and baselines on monocular depth estimation on NYUd. We freeze the backbone and perform linear probing on top of one (lin. 1) or four (lin. 4) layers.
Model Backbone lin. 1 lin. 4
RMSE $\downarrow$ RMSE $\downarrow$
ViT (Dosovitskiy et al., 2021) ViT-B/16 1.118 1.117
SAM (Kirillov et al., 2023) ViT-B/16 0.678 0.652
Theia (Shang et al., 2024) ViT-B/16 0.644 0.629
OpenCLIP (Cherti et al., 2023) ViT-G/14 0.541 0.510
MAE (He et al., 2022) ViT-H/14 0.517 0.483
DINO (Caron et al., 2021) ViT-B/8 0.555 0.539
DINOv2 (Teacher upper bound) ViT-B/14 0.399 0.362
DINOv2 (Our reproduced) ViT-B/14 0.406 0.366
[…]
SAK (Distilled from 3 teachers) ViT-B/16 0.450 0.436
Vision-language learning. To evaluate the performance of multi-modal learning, we follow the setup of RADIO (Ranzinger et al., 2024b) and integrate SAK into LLaVA-1.5 (Liu et al., 2024b; a). We freeze the Teacher-Agnostic Stem and Teacher-Specific Adapter Path, and fine-tune only the MoR Router. We evaluate three VQA tasks – GQA, POPE, and VQAv2 – at resolutions of 432 and 512, consistent with RADIO. We also include DINOv2-L, CLIP-L, and SAM-H for comparison.
Table 23 shows that while SAK is trained on significantly less data, and with smaller VFM teachers, it achieves comparable performance to the teachers and RADIO on GQA and POPE. The performance gap on VQAv2 is largely attributable to the simplicity of our distillation setup, which prioritizes training efficiency. Regarding dataset scale, as listed in Table 1 and Table 24, all VFM teachers and RADIO use large-scale datasets, e.g., DINOv2 uses 142M images, CLIP uses 400M, RADIO uses 1.4B. In contrast, our SAK distillation relies on ImageNet-1k, which has only 1.2M images. This limited dataset diversity could significantly impact performance on VQA tasks, which require extensive training data.
Table 23: Comparison with VFMs and RADIO on vision-language learning, based on LLaVA-1.5 (Liu et al., 2024a) with GQA, POPE, and VQAv2 datasets.
Model Backbone Resolution GQA POPE VQAv2
DINOv2 (Oquab et al., 2024) ViT-L/14 336 62.11 87.72 76.42
CLIP (Radford et al., 2021) ViT-L/14 336 62.20 86.09 78.49
SAM (Kirillov et al., 2023) ViT-H/16 1024 49.92 81.76 57.65
[…]
512 62.70 86.59 78.03
ViT-L/16 432 62.89 86.13 79.44
512 63.58 86.66 80.04
[…]
512 60.75 85.84 74.10
ViT-L/16 432 62.01 86.03 75.31
512 62.32 86.75 75.48
Meanwhile, RADIO employs larger VFM teachers, including DFN CLIP-H, SigLIP-SO400M, DINOv2-g, and SAM-H, which result in 2.5 times the parameters and forwarding costs of SAK with teachers based on ViT-L, shown in Table D.1. RADIO also uses a larger resolution of 1,024 during its distillation (in LLaVA experiments, the image resolution is 432 or 512), enhancing image information extraction but at a substantial increase in forwarding and backwarding costs. Considering these factors, it is reasonable that SAK cannot outperform RADIO in VLM applications. We believe that with access to larger-scale datasets and more powerful teachers during distillation, SAK could achieve competitive or superior results.
Appendix E Additional Related Work
E.1 Knowledge Distillation
@@
Knowledge Distillation (KD) is a model compression technique where a smaller student model learns from a larger teacher model, typically using the teacher’s output logits as soft targets to guide training (Buciluǎ et al., 2006; Hinton et al., 2014; Ba & Caruana, 2014; Beyer et al., 2022). To improve knowledge transfer, some studies have proposed using intermediate representations (Romero et al., 2015; Ahn et al., 2019; Heo et al., 2019; Zagoruyko & Komodakis, 2017). Multi-teacher knowledge distillation has also been explored, where knowledge is transferred from multiple teacher models simultaneously or progressively to improve the student’s generalization (Sau & Balasubramanian, 2016; You et al., 2017; Fukuda et al., 2017; Wen et al., 2024; Cao et al., 2023; Roth et al., 2024). Another branch of work, known as knowledge amalgamation, combines knowledge from multiple teachers tackling different tasks into a student that learns the union of all tasks (Shen et al., 2019a; b; Luo et al., 2019; Ye et al., 2019b; Vongkulbhisal et al., 2019; Thadajarassiri et al., 2021; 2023). Such methods mainly unify classification tasks from multiple domains. In contrast, our approach addresses multiple heterogeneous vision tasks (e.g., semantic segmentation, surface normal estimation) simultaneously.
E.2 Vision Foundation Models